# Cascading from the winds to the disk: the universality of supernovae-driven turbulence in different galactic interstellar mediums

### v1

@@ Abstract
Star-forming galaxies are in a state of turbulence, with one of the principle components of the turbulence sourced by the constant injection of momentum from supernovae (SNe) explosions. Utilizing high-resolution stratified, gravito-hydrodynamical models of SNe-driven turbulence with interstellar medium (ISM) cooling and heating, we explore how SNe-driven turbulence changes across different galactic conditions, parameterized by the galactic mass and potential, SNe-driving rate, and seeding functions. We show that even though the underlying ISM changes between starburst and Milky Way analogue models, the velocity fluctuations in the turbulence of both models, but not the kinetic energy fluctuations, can be normalized into a universal, single cascade, $u^{2}(k)\propto k^{-3/2}$, where $u$ is the velocity and $k$ is the wavemode, indicating that the structure of the turbulence is robust to significant changes in the ISM and SNe seeding. Moreover, the cascades connect smoothly from the winds into the galactic disk, pushing the outer-scale of the turbulence, $\ell_{\rm cor}$, to over $\ell_{\rm cor}\approx 6\ell_{0}$, where $\ell_{0}$ is the gaseous scale-height. By providing an analytical model for the sound speed spectrum, $c_{s}^{2}(k)$, in the weak-cooling, adiabatic limit, we show that it is the compressible turbulent modes, $u_{c}$, that control the volume-filling phase structure of the galactic disks in our models, with $c_{s}^{2}(k)\propto k^{-2}\propto u_{c}^{2}(k)$. This may indicate that galactic turbulence does not only have highly-universal features across different galaxies, but also directly sets the volume-filling hot and warm phase structure of the underlying galactic ISM through turbulent compressible modes.
@@ 1 Introduction
Padoan et al. (2016), Gent et al. (2023) and Beattie et al. (2025c) showed that there is a power law velocity and kinetic energy spectrum that resembles turbulence, albeit with deviations from the standard Kolmogorov (1941) $u^{2}(k)\propto k^{-5/3}$ model, following a shallower power law $u^{2}(k)\propto k^{-3/2}$ (Beattie et al., 2025c), even in hydrodynamical simulations, suggesting that SNe-driven turbulence has a weakened nonlinearity compared to isotropic, incompressible turbulence. Balsara et al. (2004), Padoan et al. (2016), Pan et al. (2016) and Beattie et al. (2025c) further showed that even though the energy and momentum is injected via SN blastwaves, the underlying plasma is dominated by incompressible velocity modes. Balsara et al. (2004) hypothesized that it was due to large-scale interaction between corrugated blastwaves having non-zero baroclinicity (similar to how cosmic shocks may generate magnetic fields through a Biermann battery; Kulsrud et al. 1997). However, Beattie et al. (2025c) showed conclusively that the source of the incompressible modes is from the cooling layers in SNe remnants, where the baroclinicity at the interface between hot and warm plasma phases can be many orders of magnitude larger than in the background plasma (and the other incompressible mode generation pathways). Further detailed, spectral studies of SNe-driven turbulence in different galactic environments and with different numerical prescriptions, beyond ISM phase volume-filling factors and 1-point statistics, are critical for better understanding the nature of SNe-driven turbulence in different galaxies.
@@ 2.1.2 Initial & boundary conditions
We choose a cubic simulation domain with length, $L=$1000\text{\,}\mathrm{p}\mathrm{c}$\$, allowing us to explore the large-scale ISM properties of the disk, and the disk-wind connection in the turbulence. The simulations have periodic boundary conditions on the four sides perpendicular to the disk mid-plane and outflow boundaries on the top and bottom boundaries. The domain is discretized using a regular Cartesian grid of up to $N_{\rm grid}=512$ cells for each $L$. For second-order spatial reconstruction methods such as the one used to solve our model, the numerical diffusive effects influence $\approx 10\operatorname{d}\!{x}$ (Shivakumar & Federrath, 2025; Beattie et al., 2025b), where $\operatorname{d}\!{x}=L/N_{\rm grid}$, which means that at $N_{\rm grid}=512$ we properly resolve roughly $50\text{\,}\mathrm{p}\mathrm{c}$\lesssim\ell\lesssim$1000\text{\,}\mathrm{p}\mathrm{c}$$ in our simulations, i.e., the large-scale properties of the ISM.
@@ 2.2 Supernova seeding prescriptions
The drivers of the turbulence in this study are core-collapse SNe explosions. Thermal energy, mass, and momentum are injected using the sub-grid model from Martizzi et al. (2015, 2016), which accounts for unresolved SNe evolution. The injections happen over a region of size $\ell_{\rm{inj}}=2\,\rm{dx}$ which is fixed for all resolutions (see Appendix C and Beattie et al. 2025c for spectra convergence tests and justification). The total energy, $E_{\rm{SNe}}=10^{51}\,\rm{erg}$, is partitioned based on the ambient $\rho$ and $Z$ in the local medium, into radial momentum, $p_{\rm{SNe}}$, and thermal energy, $E_{\rm{th,SNe}}$. The SNe each eject $M_{\rm{ej}}=6.8\,M_{\odot}$ new material into the ISM and continuously sweep material $M_{\rm{swept}}$ through the medium as the SNR expands. We compare two different seeding mechanisms for each galaxy type to probe their effects on the turbulence, following the “fixed” (FX) and “self-consistent” (SC) seeding prescriptions introduced in Martizzi et al. (2016), and detailed in the following section.
@@ 2.3 Statistical stationarity
Figure 2: Two-dimensional slices parallel to $\bm{\nabla}\phi$ of statistically stationary field quantities that are used throughout this study. The panels are organized by the different galactic models (MW, first two columns, and SB, last two columns) and seeding prescriptions (FX and SC suffice), with annotations shown in the top-right of each panel in the first row. The top three quantities are normalized by their respective volume-weighted root-mean-squared value. The first row is the mass density, $\rho$, the second row is the velocity magnitude, $u$, with velocity streamlines in the slice plane shown in white, the third row is the thermal sound speed, $c_{s}\propto\sqrt{P/\rho}$, and the last row is the local Mach number, $\mathcal{M}=u/c_{s}$. The MW models (first two columns) are characterized by a thick, transonic, warm disk, and a SNe explosion rate of $\gamma_{\rm SNe}/10^{-4}\,\rm{yrs}=0.1$. Whereas the SB models are characterized by a thin disk, hot, diffuse wind, and a $\gamma_{\rm SNe}$ an order of magnitude larger than the MW model.
@@ 3 ISM diversity
In Figure 2 we show two-dimensional slices through the center of the domain, parallel to $\bm{\nabla}\phi$, for each of the four galaxy and seeding models (MW and different seeding schemes in the left two columns, and SC and different seeding schemes in the right two columns). We normalize all field variables by their volume-weighted root-mean-squared value. The first row is the mass density, $\rho/\left\langle\rho^{2}\right\rangle^{1/2}$, the second row is the velocity magnitude $u/\left\langle u^{2}\right\rangle^{1/2}$ with overlaid velocity streamlines, the third row is the thermal sound speed, $c_{s}/\left\langle c_{s}^{2}\right\rangle^{1/2}$ and the final is the local $\mathcal{M}=u/c_{s}$ (for which we take the rest-frame root-mean-squared to calculate the turbulent $\mathcal{M}$ in Figure 1).
The differences between the MW and SC models are apparent in Figure 2, which will be a recurring theme in our study. The MW models feature thick, transonic, warm disks that extend throughout most of the vertical domain. In contrast, the SC models have thin, dense disks characterized by hot, subsonic winds that occupy a significant portion of the volume. This distinction is especially evident in the velocity streamlines shown in SBFX. The SNe seeding difference is particularly noticeable in the SB simulations. The FX runs show filamentary structure due to many SNe going off at the disk boundary, but none above 8080pc, ejecting dense material into the winds which cools and falls back to the disk. In the SC runs, many SNe are heating and puffing up the disk, but undergoing radiative losses in this region contributing to a less-developed wind. Overall, this dichotomy between hot, wind-dominated volume-integral statistics and warm, disk-dominated statistics between the two models makes for a robust study across two very qualitatively different ISMs.
@@ 3.1 Turbulent Mach number
For our SB model, $\mathcal{M}$ can intermittently reach $\mathcal{M}\gtrsim$. High $\mathcal{M}$ ISMs are typical of starburst galaxies (see $\mathcal{M}\approx 96$ for cold star-forming clumps in starburst M82, Sharda et al. 2018). Our model is on the low-SFR end of starbursts, more analogous to a metal-poor starburst galaxy (Kennicutt & Evans, 2012). Regardless, we find $\mathcal{M}_{\rm{SC}}=1.9\pm 0.1$ and $\mathcal{M}_{\rm{FX}}=2.0\pm 0.2$ for the volume-filling component of the ISM, comparable to our MW model and starburst galaxy ionized gas observations. Using non-thermal line broadening from the ionized gas in M82, $\left\langle u^{2}\right\rangle^{1/2}\approx 23-95\,\rm{km\,s}^{-1}$ (where we have included a factor $\sqrt{3}$, assuming isotropy, and converted the FWHM measurements into $\left\langle u^{2}\right\rangle^{1/2}=\sigma_{u}$; Westmoquette et al., 2009), and assuming a standard WIM thermal velocity $c_{s}\approx 10\,\rm{km\,s}^{-1}$, this gives $\mathcal{M}\approx 2-10$ in the WIM for M82, with a lower bound that is broadly consistent with our simulations. This shows, regardless of both the $\gamma_{\rm SNe}$ and the parameterization of the potential, the volume-integral $\mathcal{M}$ remains relatively constant across the galaxies, even when the $\gamma_{\rm SNe}$ increases by an order of magnitude, indicative of a strong covariance in $c_{s}$ and $u$ fluctuations.
@@ 3.2 Scale height & galaxy profiles
We find that $\ell_{0}$ $N_{\rm{grid}}=512^{3}$ is similar across seeding schemes for each model, with $\ell_{0,\rm{SC}}\approx 126\,\rm{pc}$ and $\ell_{0,\rm{FX}}\approx 121\,\rm{pc}$ for the MW model (in agreement with other works for the vertical gaseous disk, see Vijayakumar et al., 2025). For SB, we find $\ell_{0,\rm{SC}}\approx 43\,\rm{pc}$ and $\ell_{0,\rm{FX}}\approx 44\,\rm{pc}$, which is in agreement with observational constraints derived for gaseous disks in starbursting regions (e.g., Elmegreen et al., 2025), and radio observations of the starburst galaxy M82, (Adebahr et al., 2013, $\approx 20\rm{pc}$;). The disk is slightly thicker for the SC runs, because more SNe are seeded compactly in the midplane, contributing to pressure support and expanding the disk. As expected, the disk of the SB galaxy is considerably thinner than that of MW and incredibly wind-dominated, due to the deeper gravitational potential and higher $\gamma_{\rm{SNe}}$. This is also shown by the order-of-magnitude difference in velocity and sound speed fluctuations between MW and SB, plotted in green and purple, demonstrating that the SB host a turbulent, hot wind.
@@ 3.3 Mass density distribution functions
Figure 6: Time-averaged, two-dimensional pressure-density (top left, middle row) and temperature-density (top right and bottom row) probability distribution functions (or phase plots) for each model, with our fiducial MWFX model enlarged on the top row as a guide to key features. This illustrates the presence of a multiphase ISM with a volume-filling warm phase of $T\lesssim 10^{4}\,\rm{K}$ (WNM and WIM) and a hot $T\approx 10^{6}\,\rm{K}$ (HIM) phase. $\rho$ and $P$ have been normalized by their average midplane values, presented in Table 2. In the top panels, example adiabats $P\propto\rho^{\gamma}$ and isotherms $P\propto\rho$ are plotted, and the hot wind region (above $10^{4}\,\rm{K}$) and disk midplane (below $10^{4}\,\rm{K}$) are highlighted on the right. All cooling to $T\lesssim 10^{4}\,\rm{K}$ is achieved via adiabatic cooling, illustrated the blue, dashed adiabat, via expansion of the galaxy disk and SNR in the disk. The distributions show the varied phase structure of the ISM across seeding schemes and galaxy models.
@@ 3.4 Interstellar medium phases
For all models, there is a clear build-up of probability density below the $T=10^{4}\,\rm{K}$ isotherm around the disk midplane which lies (on average) at T=3, [^0] T=3,\!000-$3,\!500\,\rm{K}$ (see Table 2). The region above $T=10^{4}\,\rm{K}$ is primarily the warm-hot wind and outflow region from SNe feedback. We have plotted an example temperature map for the SBSC model in Figure 5, masked to show material at $T>10^{4}\,\rm{K}$ (primarily winds) and $T<10^{4}\,\rm{K}$ (mostly disk), to verify the strong temperature boundaries present between the winds and disk. As mentioned in Section 2.1.3, our cooling function mostly truncates at $T=10^{4}\,\rm{K}$, but material can still cool below this from other mechanisms, most efficiently via adiabatic expansion of the galaxy disc and SNe therein (see Equation 27). In other words, nearly all the material below $T=10^{4}\,\rm{K}$ follows an adiabat, and almost all this material is in the disk. This adiabatic trajectory, $P\propto\rho^{\gamma}$ (where $\gamma=5/3$ for monoatomic gas), is most directly visible in the phase space path from $T=10^{4}\,\rm{K}$ to $T=10^{2}\,\rm{K}$, which is plotted as a blue dashed line on the top-left plot. We also see faint adiabatic cooling tracks in the WIM and HIM, due to the expansion of SNR in the diffuse winds.
As mentioned previously, the disk of the MW model has a much higher volume-filling factor than the SB model, due to lower $\rho_{0}$ and shallower $\bm{\nabla}\phi$. Alternatively, the SB model has a higher volume-filling factor in the winds, with a significantly higher $\gamma_{\rm SNe}$and deeper $\bm{\nabla}\phi$, developing significant outflows. For both the SC simulations, and particularly for the SBSC, simulation, we see a larger and smoother spread of phase-space in the disk region and below, and therefore a higher volume of gas cooling to below $T\approx 10^{3}\,\rm{K}$ compared to SBFX, likely due to the fact that there are more SNe detonating in the dense midplane in this simulation, which increases the presence of adiabatic expansion of SNe (see Figure 5). Furthermore, for the SBSC simulation (right, second and third row), we see a buildup of probability density on an adiabatic trajectory above $T=10^{4}\,\rm{K}$, just between $\log_{10}P/P_{0}=0$ and $\log_{10}P/P_{0}=-5$. This is due to the fact that we have SNe further into the winds than is prescribed in SBFX (see the location of the SNR in Figure 5), so the SNR expand with less ambient pressure confinement and surrounding material, persisting in the Sedov-Taylor evolution for longer.
@@
Figure 7: One-dimensional, isotropic power spectra normalized by the integrated power, as a function of wavenumber, $k$, normalized by the system scale wavenumber, $2\pi/L$, with the gaseous scale height, $k_{0}\sim\ell_{0}^{-1}$, annotated on each panel with the grey vertical line. We show MW (left column) and SB (right column), for both seeding mechanisms (dark and light colors). Row one: the velocity spectrum compensated by $k^{-5/3}$ for Kolmogorov (1941)-type turbulence. Row two: the velocity spectrum decomposed into compressible $\bm{u}_{c}$ (dashed) and incompressible $\bm{u}_{s}$ (solid) modes for each seeding mechanism (offset for clarity), compensated by $k^{-2}$ Burgers (1948)-type turbulence, and with annotations for the quasi-acoustic turbulence $k^{-3}$ (Wang et al., 2017; Beattie et al., 2025b). Row three: the kinetic energy spectrum compensated by $k^{-4/3}$ (found previously in compressible turbulence Grete et al. 2021). Row four: the $c_{s}$ spectrum, with $\bm{u}_{c}$ velocity spectra modes for each simulation plotted in dashed black and grey lines, showing that both spectra trace one another other (a natural repercussion of being in the adiabatic limit and mediated by acoustic waves, Section 4.3), following $\sim k^{-2}$ for the MW simulations. We show further relevant $k$-scaling annotations in black for each panel, which we discuss in the main text.

### v2

# Cascading from the winds to the disk: the universality of supernovae-driven turbulence in different galactic interstellar media

@@ Abstract
Star-forming galaxies are in a state of turbulence, with one of the principle components of the turbulence sourced by the constant injection of momentum from supernovae (SNe) explosions. Utilizing high-resolution stratified, gravito-hydrodynamical models of SNe-driven turbulence with interstellar medium (ISM) cooling and heating, we explore how SNe-driven turbulence changes across different galactic conditions, parameterized by the galactic mass and potential, SNe-driving rate, and seeding functions. We show that even though the underlying ISM changes between starburst and Milky Way analogue models, the velocity fluctuations in the turbulence of both models, but not the kinetic energy fluctuations, can be normalized into a universal, single cascade, $du^{2}(k)/dk\propto k^{-3/2}$, where $u$ is the velocity and $k$ is the wavemode, indicating that the structure of the turbulence is robust to significant changes in the ISM and SNe seeding. Moreover, the cascades connect smoothly from the winds into the galactic disk, pushing the outer-scale of the turbulence, $\ell_{\rm cor}$, to over $\ell_{\rm cor}\approx 6\ell_{0}$, where $\ell_{0}$ is the gaseous scale-height. By providing an analytical model for the sound speed spectrum, $dc_{s}^{2}(k)/dk$, in the weak-cooling, adiabatic limit, we show that it is the compressible turbulent modes, $u_{c}$, that control the volume-filling phase structure of the galactic disks in our models, with $dc_{s}^{2}(k)/dk\propto k^{-2}\propto du_{c}^{2}(k)/dk$. This may indicate that galactic turbulence does not only have highly-universal features across different galaxies, but also directly sets the volume-filling hot and warm phase structure of the underlying galactic ISM through turbulent compressible modes.
@@ 1 Introduction
Padoan et al. (2016), Gent et al. (2023) and Beattie et al. (2025c) showed that there is a power law velocity and kinetic energy spectrum that resembles turbulence, albeit with deviations from the standard Kolmogorov (1941) $du^{2}(k)/dk\propto k^{-5/3}$ model, following a shallower power law $du^{2}(k)/dk\propto k^{-3/2}$ (Beattie et al., 2025c), even in hydrodynamical simulations, suggesting that SNe-driven turbulence has a weakened nonlinearity compared to isotropic, incompressible turbulence. Balsara et al. (2004), Padoan et al. (2016), Pan et al. (2016) and Beattie et al. (2025c) further showed that even though the energy and momentum is injected via SN blastwaves, the underlying plasma is dominated by incompressible velocity modes. Balsara et al. (2004) hypothesized that it was due to large-scale interaction between corrugated blastwaves having non-zero baroclinicity (similar to how cosmic shocks may generate magnetic fields through a Biermann battery; Kulsrud et al. 1997). However, Beattie et al. (2025c) showed conclusively that the source of the incompressible modes is from the cooling layers in SNe remnants, where the baroclinicity at the interface between hot and warm plasma phases can be many orders of magnitude larger than in the background plasma (and the other incompressible mode generation pathways). Further detailed, spectral studies of SNe-driven turbulence in different galactic environments and with different numerical prescriptions, beyond ISM phase volume-filling factors and 1-point statistics, are critical for better understanding the nature of SNe-driven turbulence in different galaxies.
@@ 2.1.2 Initial & boundary conditions
We choose a cubic simulation domain with length, $L=$1000\text{\,}\mathrm{p}\mathrm{c}$\$, allowing us to explore the large-scale ISM properties of the disk, and the disk-wind connection in the turbulence. The simulations have periodic boundary conditions on the four sides perpendicular to the disk mid-plane and outflow boundaries on the top and bottom boundaries. The outflow boundaries suppress any inflowing material from the ghost cells. The domain is discretized using a regular Cartesian grid of up to $N_{\rm grid}=512$ cells for each $L$. For second-order spatial reconstruction methods such as the one used to solve our model, the numerical diffusive effects influence $\approx 10\operatorname{d}\!{x}$ (Shivakumar & Federrath, 2025; Beattie et al., 2025b), where $\operatorname{d}\!{x}=L/N_{\rm grid}$, which means that at $N_{\rm grid}=512$ we properly resolve roughly $50\text{\,}\mathrm{p}\mathrm{c}$\lesssim\ell\lesssim$1000\text{\,}\mathrm{p}\mathrm{c}$$ in our simulations, i.e., the large-scale properties of the ISM.
@@ 2.2 Supernova seeding prescriptions
The drivers of the turbulence in this study are core-collapse SNe explosions. Thermal energy, mass, and momentum are injected using the sub-grid model from Martizzi et al. (2015, 2016), which accounts for unresolved SNe evolution. The injections happen over a region of size $\ell_{\rm{inj}}=2\,\rm{dx}$ which is fixed for all resolutions (see Appendix D and Beattie et al. 2025c for spectra convergence tests and justification). The total energy, $E_{\rm{SNe}}=10^{51}\,\rm{erg}$, is partitioned based on the ambient $\rho$ and $Z$ in the local medium, into radial momentum, $p_{\rm{SNe}}$, and thermal energy, $E_{\rm{th,SNe}}$. The SNe each eject $M_{\rm{ej}}=6.8\,M_{\odot}$ new material into the ISM and continuously sweep material $M_{\rm{swept}}$ through the medium as the SNR expands. We compare two different seeding mechanisms for each galaxy type to probe their effects on the turbulence, following the “fixed” (FX) and “self-consistent” (SC) seeding prescriptions introduced in Martizzi et al. (2016), and detailed in the following section.
@@ 2.3 Statistical stationarity
Figure 2: Two-dimensional slices parallel to $\bm{\nabla}\phi$ of statistically stationary field quantities that are used throughout this study. The panels are organized by the different galactic models (MW, first two columns, and SB, last two columns) and seeding prescriptions (FX and SC suffixes), with annotations shown in the top-right of each panel in the first row. The top three quantities are normalized by their respective volume-weighted root-mean-squared value. The first row is the mass density, $\rho$, the second row is the velocity magnitude, $u$, with velocity streamlines in the slice plane shown in white, the third row is the thermal sound speed, $c_{s}\propto\sqrt{P/\rho}$, and the last row is the local Mach number, $\mathcal{M}=u/c_{s}$. The MW models (first two columns) are characterized by a thick, transonic, warm disk, and a SNe explosion rate of $\gamma_{\rm SNe}/10^{-4}\,\rm{yrs}=0.1$. Whereas the SB models are characterized by a thin disk, hot, diffuse wind, and a $\gamma_{\rm SNe}$ an order of magnitude larger than the MW model.
@@ 3 ISM diversity
In Figure 2 we show two-dimensional slices through the center of the domain, parallel to $\bm{\nabla}\phi$, for each of the four galaxy and seeding models (MW and different seeding schemes in the left two columns, and SB and different seeding schemes in the right two columns). We normalize all field variables by their volume-weighted root-mean-squared value. The first row is the mass density, $\rho/\left\langle\rho^{2}\right\rangle^{1/2}$, the second row is the velocity magnitude $u/\left\langle u^{2}\right\rangle^{1/2}$ with overlaid velocity streamlines, the third row is the thermal sound speed, $c_{s}/\left\langle c_{s}^{2}\right\rangle^{1/2}$ and the final is the local $\mathcal{M}=u/c_{s}$ (for which we take the rest-frame root-mean-squared to calculate the turbulent $\mathcal{M}$ in Figure 1).
The differences between the MW and SB models are apparent in Figure 2, which will be a recurring theme in our study. The MW models feature thick, transonic, warm disks that extend throughout most of the vertical domain. In contrast, the SB models have thin, dense disks characterized by hot, subsonic winds that occupy a significant portion of the volume. This distinction is especially evident in the velocity streamlines shown in SBFX. The SNe seeding difference is particularly noticeable in the SB simulations. The FX runs show filamentary structure due to many SNe going off at the disk boundary, but none above 8080pc, ejecting dense material into the winds which cools and falls back to the disk. In the SC runs, many SNe are heating and puffing up the disk, but undergoing radiative losses in this region contributing to a less-developed wind. Overall, this dichotomy between hot, wind-dominated volume-integral statistics and warm, disk-dominated statistics between the two models makes for a robust study across two very qualitatively different ISMs.
@@ 3.1 Turbulent Mach number
For our SB model, $\mathcal{M}$ can intermittently reach $\mathcal{M}\gtrsim 10$. High $\mathcal{M}$ ISMs are typical of starburst galaxies (see $\mathcal{M}\approx 96$ for cold star-forming clumps in starburst M82, Sharda et al. 2018). Our model is on the low-SFR end of starbursts, more analogous to a metal-poor starburst galaxy (Kennicutt & Evans, 2012). Regardless, we find $\mathcal{M}_{\rm{SC}}=1.9\pm 0.1$ and $\mathcal{M}_{\rm{FX}}=2.0\pm 0.2$ for the volume-filling component of the ISM, comparable to our MW model and starburst galaxy ionized gas observations. Using non-thermal line broadening from the ionized gas in M82, $\left\langle u^{2}\right\rangle^{1/2}\approx 23-95\,\rm{km\,s}^{-1}$ (where we have included a factor $\sqrt{3}$, assuming isotropy, and converted the FWHM measurements into $\left\langle u^{2}\right\rangle^{1/2}=\sigma_{u}$; Westmoquette et al., 2009), and assuming a standard WIM thermal velocity $c_{s}\approx 10\,\rm{km\,s}^{-1}$, this gives $\mathcal{M}\approx 2-10$ in the WIM for M82, with a lower bound that is broadly consistent with our simulations. This shows, regardless of both the $\gamma_{\rm SNe}$ and the parameterization of the potential, the volume-integral $\mathcal{M}$ remains relatively constant across the galaxies, even when the $\gamma_{\rm SNe}$ increases by an order of magnitude, indicative of a strong covariance in $c_{s}$ and $u$ fluctuations.
@@ 3.2 Scale height & galaxy profiles
We find in Figure 3 that $\ell_{0}$ is similar across seeding schemes for each model, with $\ell_{0,\rm{SC}}\approx 126\,\rm{pc}$ and $\ell_{0,\rm{FX}}\approx 121\,\rm{pc}$ for the MW model (in agreement with other works for the vertical gaseous disk, see Vijayakumar et al., 2025). For SB, we find $\ell_{0,\rm{SC}}\approx 43\,\rm{pc}$ and $\ell_{0,\rm{FX}}\approx 44\,\rm{pc}$, which is in agreement with observational constraints derived for gaseous disks in starbursting regions (e.g., Elmegreen et al., 2025), and radio observations of the starburst galaxy M82, (Adebahr et al., 2013, $\approx 20\rm{pc}$;). The disk is slightly thicker for the SC runs, because more SNe are seeded compactly in the midplane, contributing to pressure support and expanding the disk. As expected, the disk of the SB galaxy is considerably thinner than that of MW and incredibly wind-dominated, due to the deeper gravitational potential and higher $\gamma_{\rm{SNe}}$. This is also shown by the order-of-magnitude difference in velocity and sound speed fluctuations between MW and SB, plotted in green and purple, demonstrating that the SB host a turbulent, hot wind.
@@ 3.3 Mass density distribution functions
Figure 6: Time-averaged, two-dimensional pressure-density (top left, middle row) and temperature-density (top right, bottom row) probability distribution functions (or phase plots) for each model, normalized by average midplane values in steady state. Our fiducial MWFX model is enlarged on the top row as a guide to key features. This illustrates the presence of a multiphase ISM with a volume-filling warm phase of $T\lesssim 10^{4}\,\rm{K}$ (WNM and WIM) and a hot $T\approx 10^{6}\,\rm{K}$ (HIM) phase. $\rho$ and $P$ have been normalized by their average midplane values, presented in Table 2. In the top panels, example adiabats $P\propto\rho^{\gamma}$ and isotherms $P\propto\rho$ are plotted, and the hot wind region (above $10^{4}\,\rm{K}$) and disk midplane (below $10^{4}\,\rm{K}$) are highlighted on the right. All cooling to $T\lesssim 10^{4}\,\rm{K}$ is achieved via adiabatic cooling, illustrated by the example blue, dashed adiabat, via expansion of the galaxy disk and SNR in the disk. The distributions show the varied phase structure of the ISM across seeding schemes and galaxy models.
@@ 3.4 Interstellar medium phases
For all models, there is a clear build-up of flat probability density below the $T=10^{4}\,\rm{K}$ isotherm around the disk midplane which lies (on average) at T=3, [^0] T=3,\!000-$3,\!500\,\rm{K}$ (see Table 2). The region above $T=10^{4}\,\rm{K}$ is primarily the warm-hot wind and outflow region from SNe feedback. We have plotted an example temperature map for the SBSC model in Figure 5, masked to show material at $T>10^{4}\,\rm{K}$ (primarily winds) and $T<10^{4}\,\rm{K}$ (mostly disk), to verify the strong temperature boundaries present between the winds and disk. As mentioned in Section 2.1.3, our cooling function mostly truncates at $T=10^{4}\,\rm{K}$, but material can still cool below this from other mechanisms, most efficiently via adiabatic expansion of the galaxy disc and SNe therein (see Equation 27). In other words, nearly all the material below $T=10^{4}\,\rm{K}$ follows an adiabat, and almost all this material is in the disk. This adiabatic trajectory, $P\propto\rho^{\gamma}$ (where $\gamma=5/3$ for monoatomic gas), is most directly visible in the phase space paths from $T=10^{4}\,\rm{K}$ to $T=10^{2}\,\rm{K}$, one of which is plotted as a blue dashed line on the top-left plot. Any deviation from these adiabatic trajectories is likely due to the non-zero cooling function having small effects in the lower temperature phases, detailed in Section 2.1.3. We also see faint adiabatic cooling tracks in the WIM and HIM, due to the expansion of SNR in the diffuse winds.
As mentioned previously, the disk of the MW model has a much higher volume-filling factor than the SB model, due to lower $\rho_{0}$ and shallower $\bm{\nabla}\phi$. Alternatively, the SB model has a higher volume-filling factor in the winds, with a significantly higher $\gamma_{\rm SNe}$and deeper $\bm{\nabla}\phi$, developing significant outflows. For both the SC simulations, and particularly for the SBSC, simulation, we see a larger and smoother spread of phase-space in the disk region and below, and therefore a higher volume of gas cooling to below $T\approx 10^{3}\,\rm{K}$ compared to SBFX, likely due to the fact that there are more SNe detonating in the dense midplane in this simulation, which increases the presence of adiabatic expansion of SNe (see Figure 5). Furthermore, for the SBSC simulation (right, second and third row), we see a buildup of probability density on an adiabatic trajectory above $T=10^{4}\,\rm{K}$, just between $\log_{10}P/P_{0}=0$ and $\log_{10}P/P_{0}=-5$. This is due to the fact that we have SNe further into the winds than is prescribed in SBFX (see the location of the SNR in Figure 5), so the SNR expand with less ambient pressure confinement and surrounding material, persisting in the adiabatic Sedov-Taylor evolution for longer.
@@
Figure 7: One-dimensional, isotropic power spectra normalized by the integrated power, as a function of wavenumber, $k$, normalized by the system scale wavenumber, $2\pi/L$, with the gaseous scale height, $k_{0}\sim\ell_{0}^{-1}$, annotated on each panel with the grey vertical line. The shaded regions show the $1\sigma$ fluctuations of the spectra in time. We show MW (left column) and SB (right column), for both seeding mechanisms (dark and light colors). All fitted indexes for spectra are shown in Table 3. Row one: the velocity spectrum compensated by $k^{-5/3}$ for Kolmogorov (1941)-type turbulence. Row two: the velocity spectrum decomposed into compressible $\bm{u}_{c}$ (dashed) and incompressible $\bm{u}_{s}$ (solid) modes for each seeding mechanism (offset for clarity), compensated by $k^{-2}$ Burgers (1948)-type turbulence, and with annotations for the quasi-acoustic turbulence $k^{-3}$ (Wang et al., 2017; Beattie et al., 2025b). Row three: the kinetic energy spectrum compensated by $k^{-4/3}$ (found previously in compressible turbulence Grete et al. 2021). Row four: the $c_{s}$ spectrum, with $\bm{u}_{c}$ velocity spectra modes for each simulation plotted in dashed black and grey lines, showing that both spectra trace one another other (a natural repercussion of being in the adiabatic limit and mediated by acoustic waves, Section 4.3), following $\sim k^{-2}$ for the MW simulations. We show further relevant $k$-scaling annotations in black for each panel, which we discuss in the main text.